# A Spatial-statistical model to analyse historical rutting data

### 08 Jan 2024 v1

@@ A.3 Marginal Posterior Distributions of Random Effects
Figure A.3 gives the marginal posterior distributions of the standard deviation of the Gaussian observations ($\hat{\sigma}_{\varepsilon}$), the standard deviation of the random yearly effect ($\hat{\sigma}_{\gamma}$), the range of the common spatial field $\hat{\rho}$, and the standard deviation of the common spatial field $\hat{\sigma}_{\omega(s)}$. The solid lines represent the spatial models and the dotted lines represent the non-spatial models. Narrower densities suggest smaller estimated uncertainty. We find that some of the uncertainty of the Gaussian observations $(\hat{\sigma}_{\varepsilon})$ and random yearly effect $(\hat{\sigma}_{\gamma})$ go into the spatial effect $\{\hat{\sigma}_{\omega(s)}\}$, and this can be interpreted as some of the unobserved variability (e.g., weather, temperature) being explained by the spatial component; see Table 3 and Figure A.3 (a) and (b). This is confirmed by the smaller estimates of the ‘noise’ term $(\hat{\sigma}_{\varepsilon})$ compared with the non-spatial model (Table 4).
We examine the effects of the range of the common spatial field ($\hat{\rho}$), and the standard deviation of the common spatial field $\{\hat{\sigma}_{\omega(s)}\}$. As more explanatory variables are added to model (e.g., rut depth), the variability of the common spatial field $\{\hat{\sigma}_{\omega(s)}\}$ decreases; see Figure A.3, graph (d).

### 14 Feb 2024 v2

@@ A.3 Marginal Posterior Distributions of Random Effects
Figure A.3 gives the marginal posterior distributions of the standard deviation of the Gaussian observations ($\hat{\sigma}_{\varepsilon}$), the standard deviation of the random yearly effect ($\hat{\sigma}_{\gamma}$), the range of the common spatial field $\hat{\rho}$, and the standard deviation of the common spatial field $\hat{\sigma}_{\omega(s)}$. The solid lines represent the spatial models and the dotted lines represent the non-spatial models. Narrower densities suggest smaller estimated uncertainty. We find that some of the uncertainty of the Gaussian observations $(\hat{\sigma}_{\varepsilon})$ and random yearly effect $(\hat{\sigma}_{\gamma})$ go into the spatial effect $\{\hat{\sigma}_{\omega(s)}\}$, and this can be interpreted as some of the unobserved variability (e.g., weather, temperature) being explained by the spatial rutting component; see Table 3 and Figure A.3 (a) and (b). This is confirmed by the smaller estimates of the ‘noise’ term $(\hat{\sigma}_{\varepsilon})$ compared with the non-spatial model (Table 4).
We examine the effects of the range of the common spatial field ($\hat{\rho}$), and the standard deviation of the common spatial field $\{\hat{\sigma}_{\omega(s)}\}$. As more explanatory variables are added to model (e.g., rut depth), the variability of the common spatial field $\{\hat{\sigma}_{\omega(s)}\}$ decreases; see Table 4 and Figure A.3, graph (d).
@@
A.4 Area with high rutting rates
Figure A.4 shows the location where rutting exceeds $7$ mm when the worse values of the explanatory variables are used to explore Model 1 (see Section 4.4)
(a) Short section between EV14 S6D1 m6802-m6882 F1 with high rutting.
(b) Location EV14 S6D1 m6881 F1 in year 2020
(c) Location EV14 S6D1 m6808 F1 in year 2021
(d) Location EV14 S6D1 m6807 F1 in year 2022
Figure A.4: Graph (a): Location with high rutting rates (road segments $2100-2700$ towards Sweden). Graphs (b), (c) and (d) show the location, in years 2020-2022 (see Statens vegvesen or go to the next url: https://vegkart.atlas.vegvesen.no/#kartlag:nib/@337039,7037105,18/vegsystemreferanse:337039.29:7037149.25